OpenStudy (anonymous):

explain how to factor trinomial forms x^2 + bc +c and ax^2 + bx + c

OpenStudy (happinessbreaksbones):

hmm well I'm still new to this and I hope my answer is right.. but I believe it would be factored as: \[x ^{2} +bc + c\] example equation: \[2x^2 - 2x - 24\] the answer would be \[2(x - 4)(x + 3)\] because you have to distribute as well as FOIL.. x • x = x^2 and 2 • x^2 = 2x^2 -4 • 3 = -12 and 2 • -12 would be -24 and as for -2x in the middle.. continue to FOIL -4 • x = -4x and -4x + 3 = -1x.. -1x • 2 (on the outside) = -2x.. so \[2x^2 - 2x - 24\] would equal \[2(x - 4)(x + 3)\] I really hope this helped. Hopefully if not, someone else can help or I can further explain?